Find the equation through the point (2,4) with slope 4
Answer by anantha(86)

sol:
equation of the line passing through (x1,y1) and slope m is
y-y1=m(x-x1)
in this problem the coordinates (x1,y1)=(2,4) means x1=2 and y1=4 and slope m=4
substitute these values in the above formula
y-4=4(x-2)
y-4=4x-8
transpose L.H.S part to R.H.S
4x-8-y+4=0
4x-y-4=0
which is the requried equation passing through (2,4) and slope 4